OpenStudy (anonymous):

find the derivative of y = sqrt(3)+(x)^1/3 + 1/x???

OpenStudy (anonymous):

\[\sqrt{x}+x^{\frac{1}{3}}+\frac{1}{x}\]?

OpenStudy (anonymous):

so the problem is A) \[y=\sqrt{3+x ^{1/3}+1/x}\] or B) \[y=\sqrt{3}+x ^{1/3}+1/x\]

OpenStudy (anonymous):

or is first term \[\sqrt{3}\]?

OpenStudy (anonymous):

the b) style

OpenStudy (anonymous):

then do not be hoodwinked. \[\sqrt{3}\] is constant, its derivative is 0

OpenStudy (anonymous):

yeap derivative of a constant is zero

OpenStudy (anonymous):

satelllite ur dis ans is wrong it does not tally wid de book pls explain in detail

OpenStudy (anonymous):

oh i did not give a complete answer. sorry. i only said derivative of \[\sqrt{3}\] is 0 the others need power rule.

OpenStudy (anonymous):

(1/3)x^-2/3-1x^-2

OpenStudy (anonymous):

however, \[\frac{1}{x}\] os a very common function, so you should just remember its derivative without using power rule. the derivative of \[\frac{1}{x}\] is \[-\frac{1}{x^2}\]

OpenStudy (anonymous):

pls don't be angry at me cause i am in 11th and just a learner!!!!!!

OpenStudy (anonymous):

Guys I don't know how to write the equation here coz I am bad at latex stuffs

OpenStudy (anonymous):

k i will telll wat ans de book has

OpenStudy (anonymous):

comes up so often you do not want to compute it afresh each time. for \[x^{\frac{1}{3}}\] youi use the power rule to get \[\frac{1}{3}x^{-\frac{2}{3}}\]

OpenStudy (anonymous):

I hope tejeshwar95 got the answer

OpenStudy (anonymous):

it says - 1/2*sqrt(3/x)+1/3*1/x^(2/3) - 1/x^2

OpenStudy (anonymous):

which is the same as \[\frac{1}{2 \sqrt [3]{x^2}}\]

OpenStudy (anonymous):

can't understand a thing

OpenStudy (anonymous):

sorry lets go slow. is the first term just \[\sqrt{3}\] or is it \[\sqrt{3x}\]

OpenStudy (anonymous):

sorry for such a rude reply but i really can't get a hang of it!!!!! its sqrt(3x)

OpenStudy (anonymous):

ooooh then i apologize. i thought it was just \[\sqrt{3}\]

OpenStudy (anonymous):

its ok!!!!!!!!!!

OpenStudy (anonymous):

lets do them one at a time. you want to use the power rule, which says that the derivative of \[x^r\] is \[r\times x^{r-1}\] ok so far?

OpenStudy (anonymous):

wat is de power rule??? haven't heard bout it

OpenStudy (anonymous):

it tells you how to find the derivative of something raised to a power. for example the derivative of \[x^3\] is \[3x^2\]

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

ok the 1 which says if y = x^n then n(x)^n-1

OpenStudy (anonymous):

yups got it till here!!!!!!

OpenStudy (anonymous):

yes that is the power rule.

OpenStudy (anonymous):

so the trick is to write each of these terms in terms of exponents and then use the power rule.

OpenStudy (anonymous):

k let me try it on paper hang on for a minute!!!!!!

OpenStudy (anonymous):

now \[x^{\frac{1}{3}}\] is already written that way so that one is easy.

OpenStudy (anonymous):

i will wait.

OpenStudy (anonymous):

as far as \[\sqrt{3}\] isconcerned we remove the 3??? cause we can't differentiate it???

OpenStudy (anonymous):

this one is the confusing one, but don't be fooled. \[\sqrt{3x}=\sqrt{3} \times \sqrt{x} = \sqrt{3}\times {x^{\frac{1}{2}}}\] and a constant just stays there.

OpenStudy (anonymous):

k got it till here

OpenStudy (anonymous):

for example, the derivative of \[x^2 \] is \[2x\] and the derivative of \[\sqrt{3}x^2\] is \[2 \sqrt{3}x\] the constant just says as a multiplier, so ignore it.

OpenStudy (anonymous):

k so i ignore sqrt(3)

OpenStudy (anonymous):

so only use the power rule on the \[x^\frac{1}{2}\] part. bring out the exponent as a multiplier, and then subtract 1 from the exponent. i wait while you try it.

OpenStudy (anonymous):

but now did u take \[\sqrt{3}x ^{2}\] come into picture

OpenStudy (anonymous):

u took it as an eg???

OpenStudy (anonymous):

that was just an example to explain that the \[\sqrt{3}\] is unimportant. just a side example. not part of this problem.

OpenStudy (anonymous):

yes just eg.

OpenStudy (anonymous):

k now after trying i got 1/2*x ^-1/2

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

now convert back to radical from from exponential form.

OpenStudy (anonymous):

k correct till here sorry for being so slow

OpenStudy (anonymous):

no problems. long as you understand.

OpenStudy (anonymous):

do the same again??????

OpenStudy (anonymous):

do you know what \[x^{-\frac{1}{2}}\] is in radical form?

OpenStudy (anonymous):

if not i explain, if so just convert back.

OpenStudy (anonymous):

nope not exactly

OpenStudy (anonymous):

ok. i explain. the exponent has a minus sign, so that means take the reciprocal. for example, \[x^{-5}=\frac{1}{x^5}\]

OpenStudy (anonymous):

yups got it

OpenStudy (anonymous):

the exponent is a fraction. the denominator is 2, so that means take the square root. the numerator is 1, so raise it to the power of 1, which is like doing nothing. so \[x^{-\frac{1}{2}}=\frac{1}{\sqrt{x}}\]

OpenStudy (anonymous):

therefore \[\frac{1}{2}x^{-\frac{1}{2}}=\frac{1}{2\sqrt{x}}\]

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

so we put \[\sqrt{x^3}\] or \[\sqrt{x}\]

OpenStudy (anonymous):

in the denominator!!!!!!

OpenStudy (anonymous):

just \[\sqrt{x}\] in the denominator. but also a 2 in the denominator because you are multiplying by 1/2

OpenStudy (anonymous):

and the numerator is this case is \[\sqrt{3}\] because that constant is still there.

OpenStudy (anonymous):

can we have a voice chat???? dis is becoming a pain!!!!!!

OpenStudy (anonymous):

final answer: \[\frac{\sqrt{3}}{2\sqrt{x}}\]

OpenStudy (anonymous):

don't know how to voice chat. do you?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

voltage drops are happening i might not reply in between

OpenStudy (anonymous):

ok. the reason this problem is a pain for you is that you have to do three things: 1)convert to exponential form 2) use the power rule 3) convert back to radical form

OpenStudy (anonymous):

if u can hang around then give me some time i will just go through dese rules!!!!!!!!

OpenStudy (anonymous):

but you probably know what \[7\times 8\] is because you have it memorized. since you are taking calculus, and since \[\sqrt{x}\] is such a common function, you should probably memorize its derivative, which is \[\frac{1}{2\sqrt{x}}\] that way you never have to do this again!

OpenStudy (anonymous):

saves you the three steps of writing in exponential form, using the power rule, and converting back. if you remember it then for homework or on a test you just write it.

OpenStudy (anonymous):

k learnt

OpenStudy (anonymous):

and derived

OpenStudy (anonymous):

and another very common function is \[f(x)=\frac{1}{x}\]

OpenStudy (anonymous):

its derivative is \[f'(x)=-\frac{1}{x^2}\]

OpenStudy (anonymous):

k got it

OpenStudy (anonymous):

but could not derive it!!!!!!!!

OpenStudy (anonymous):

note the "-" sign. you can do this using the power rule as well, but it never changes. \[\frac{1}{x}=x^{-1}\] power rule gives \[-1\times x^{-2}=-\frac{1}{x^2}\]

OpenStudy (anonymous):

when do u come online tomorrow i am tired i need to relax!!!!!!!!!!

OpenStudy (anonymous):

or give me ur no if u live in delhi den i can talk 2 u over de phone!!!!!!!

OpenStudy (anonymous):

probably in the morning if you are here then. review power rule and of course exponents (because that is what it uses) in the mean time. good luck!

OpenStudy (anonymous):

no i am in us.

OpenStudy (anonymous):

k so wat's de time dere now????????

OpenStudy (anonymous):

thanx for ur help and the pains u took but if u cud come online at the same time as u came 2day den it wud be gr8

OpenStudy (anonymous):

ok i will try. look for me around this time or a little earlier.

OpenStudy (anonymous):

like wat's de time in US?? then i can guess wen 2 come online!!!!!!

OpenStudy (anonymous):

or can u hang around for a while ny the time i play a bit of COD

OpenStudy (anonymous):

k luks like i got it!!!!!!!!!!!

OpenStudy (anonymous):

ok i will be here for a while. have some work to do.

OpenStudy (anonymous):

i got the simplified form but ain't getin de full ans!!!!!!!!!!!

OpenStudy (anonymous):

to part 1?

OpenStudy (anonymous):

I got till see attachment

OpenStudy (anonymous):

avoid the torn part look at the 1 written below!!!!!!!

OpenStudy (anonymous):

looks good to me. of course you still have actually subtract the exponents and convert back to radical form.

OpenStudy (anonymous):

how do i do dat pls tell!!!!!!

OpenStudy (anonymous):

lets do the middle one. \[\frac{1}{3}x^{\frac{1}{3}-1}=\frac{1}{3}x^{-\frac{2}{3}}\]

OpenStudy (anonymous):

so far so good?